# Practical, Utilitarian Algorithm Configuration

### v1

@@ Abstract
Utilitarian algorithm configuration identifies a parameter setting for a given algorithm that maximizes a user’s utility. Utility functions offer a theoretically well grounded approach to optimizing decision-making under uncertainty and are flexible enough to capture a user’s preferences over algorithm runtimes (e.g., they can describe a sharp cutoff after which a solution is no longer required, a per-hour cost for compute, or diminishing returns from algorithms that take longer to run). COUP is a recently-introduced utilitarian algorithm configuration procedure which was designed mainly to offer strong theoretical guarantees about the quality of the configuration it returns, with less attention paid to its practical performance. This paper closes that gap, bringing theoretically-grounded, utilitarian algorithm configuration to the point where it is competitive with widely used, heuristic configuration procedures that offer no performance guarantees. We present a series of improvements to COUP that improve its empirical performance without degrading its theoretical guarantees and demonstrate their benefit experimentally. Using a case study, we also illustrate ways of exploring the robustness of a given solution to the algorithm selection problem to variations in the utility function.
Codehttps://github.com/drgrhm/coup-plus
@@ 1 Introduction
Algorithm configuration is the process of finding a complete setting of an algorithm’s parameters that lead it to perform well on a given set of input instances. Often, the algorithm of interest is a solver for a hard computational problem such as Boolean Satisfiability (SAT). Since SAT is NP-complete, we are unlikely to be able to find a single solver or configuration that performs well on all possible inputs. In practice, however, we are never interested in solving all possible inputs. Instead, we tend to be interested in solving input instances drawn from a given distribution. Algorithm configuration is thus a bandit problem (Lattimore and Szepesvári 2020) where each configuration is an arm, and running a configuration on an input instance means drawing a sample from that arm’s underlying distribution. Unlike the standard bandit setting, in which there is a unit cost associated with pulling an arm, in algorithm configuration the cost is the time it takes for the algorithm to solve the input. We can also choose whether each such evaluation should be censored (“capped”) at some maximum runtime. The objective in algorithm configuration is not regret minimization, but rather (approximate) best arm identification: we are willing to spend time up front searching for a good configuration.
@@
Algorithm configuration procedures have traditionally aimed to minimize the expected runtime of the returned configuration. The most practical such methods make these decisions according to carefully chosen heuristics (Birattari et al. 2002; Hutter et al. 2009; Ansótegui, Sellmann, and Tierney 2009; Hutter, Hoos, and Leyton-Brown 2011; López-Ibáñez et al. 2016; Lindauer et al. 2022). A second line of work has identified (considerably less practical) algorithms that offer worst-case guarantees about the runtime required to identify approximately optimal solutions (Kleinberg, Leyton-Brown, and Lucier 2017; Weisz, Gyorgy, and Szepesvári 2018; Kleinberg et al. 2019; Weisz, Gyorgy, and Szepesvári 2019; Weisz et al. 2020). For the purposes of this paper, we define an optimality guarantee as depending on three parameters: $\epsilon$, $\gamma$, and $\delta$. With failure probability $\delta$, we want to prove both that the returned configuration is within an additive factor of $\epsilon$ from the best configuration sampled, and that the probability of sampling a configuration that is better than the best one sampled so far is at most $\gamma$.
@@
It has recently been argued that algorithm configuration should rest on the firmer decision-theoretic foundation of maximizing a utility function over algorithm runtime (Graham, Leyton-Brown, and Roughgarden 2023a). That is, let $u$ be a function that maps runtimes to utility values in $[0,1]$, encoding the user’s preferences about runtimes of different lengths; we should then maximize $u$’s expected value. A recent line of work has proposed procedures that offer worst-case guarantees about the runtime required to approximately optimize a utilitarian objective (Graham, Leyton-Brown, and Roughgarden 2023b; Brandt et al. 2023). The most recent such procedure is COUP (Continuous, Optimistic Utilitarian Procrastination; Graham and Leyton-Brown 2024), which was shown empirically to perform favorably with respect to other configuration procedures that offer (in some cases, somewhat incomparable) optimality guarantees: Successive Halving (Jamieson and Talwalkar 2016), Hyperband (Li et al. 2018), AC-Band (Brandt et al. 2023), ImpatientCapsAndRuns (Weisz et al. 2020) and Structured Procrastination with Confidence (Kleinberg et al. 2019). Pseudocode for COUP is included in Appendix A, but in essence it is an instantiation of the UCB algorithm (Lai and Robbins 1985; Lai 1987), coupled with a captime exploration technique and a condition for expanding the set of configurations under consideration. For each configuration $i$ COUP maintains upper and lower confidence bounds $UCB_{i}$ and $LCB_{i}$. In each iteration it chooses the configuration $i^{\prime}=\arg\max_{i}UCB_{i}$ and performs a run of it. In this way it works to improve $\epsilon^{*}=\max_{i}UCB_{i}-\max_{i}LCB_{i}$, which is the best optimality guarantee it can make. If COUP detects that the captime it is using is too small, it will double it. If COUP detects that its set of configurations is too small, it will add more. COUP is anytime in its guarantees with respect to both $\epsilon$ and $\gamma$: as it runs it makes tighter and tighter optimality guarantees with respect to larger and larger sets of configurations. COUP is also adaptive to the inputs it is given: it provably makes better guarantees in less time when given easier inputs.
@@
This paper begins in Section 2 by describing a series of improvements to COUP that improve its empirical performance without compromising its theoretical guarantees. First, we refine the confidence bounds used by COUP, making the optimization process more efficient. Second, we incorporate a different version of the UCB algorithm that specifically targets the best arm identification problem. Third, we suggest a new condition for when COUP should add new configurations, which frees the user from having to provide a set of parameters ahead of time, instead adapting to the inputs it sees. And fourth, we make COUP model-based, using a boosted forest of regression trees to help COUP find new configurations. Next, in Section 3 we show for the first time that an algorithm configuration method offering theoretical guarantees (our modified version of COUP) robustly achieves performance roughly equivalent to that of SMAC (Hutter, Hoos, and Leyton-Brown 2011; Lindauer et al. 2022), a heuristic procedure widely seen as the empirical state of the art. Finally, in Section 4, we address a shortcoming of the utilitarian framework: that it is not always easy for the end user to choose the exact utility function to optimize. Via a case study on data from one of the most prominent international evaluations of heuristic algorithms, the SAT Competition, we illustrate how runtime data can be used to reason about the utility functions for which a given algorithm would be considered optimal.
@@ 2.1 Improved Confidence Bounds
A key technical tool used in COUP is Hoeffding’s Inequality (Hoeffding 1963). This concentration inequality asserts that for bounded random variables, the probability that the observed sample mean is far from the true population mean is exponentially small in the number of samples taken. However, Hoeffding’s Inequality involves a key approximation, which is tight when the true mean is close to $1/2$ (in the case of random variables bounded on [0,1]), but grows worse when the true mean is close to either 0 or 1—a situation encountered often in utilitarian algorithm configuration. We can do away with this approximation by appealing directly to the “Chernoff–Hoeffding Lemma”, and solving numerically for the confidence bounds, which are then guaranteed to be tighter than a naive application of Hoeffding’s Inequality; for a detailed exposition see Lattimore and Szepesvári (2020), Chapter 10. The confidence bounds come from solving an optimization problem that bounds the KL-Divergence between the observed sample mean and the bounding quantity. We thus refer to them as “KL bounds”.
COUP maintains a captime $\kappa$, which is the maximum time it will spend each time it runs a configuration. To derive upper and lower confidence bounds for a configuration $i$’s mean utility, COUP needs to estimate its capped mean runtime $U_{i}=\mathbb{E}_{t_{i}}[u(\min(t_{i},\kappa))]$ and its probability of not capping $F_{i}=\Pr_{t_{i}}(t_{i}\leq\kappa)$. Both of these estimates can be improved by switching from Hoeffding bounds to KL bounds. Define $d(p,q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$. Suppose $n$ is the current number of configurations, $m$ is the number of samples we have taken for configuration $i$, and $\kappa$ is the captime we have taken them at. If $\widehat{U}_{i}$ is the empirical capped average runtime, and $\widehat{F}_{i}$ is the empirical fraction of completed runs, and $a=\frac{1}{m}\ln\frac{36n^{2}m^{2}(\log(\kappa+1)^{2}}{\delta}$ then
@@
are confidence bounds that hold for all configurations at all times during COUP’s execution with probability at least $1-\delta$ (detailed derivations can be found in Appendix B). This alteration essentially makes COUP an implementation of the KL-UCB algorithm (Cappé et al. 2013) for bandit problems. In Figure 2 (Left) in Section 3 we show that COUP performs better when using these refined bounds.
@@ 2.2 Lower Upper Confidence Bound
The UCB algorithm is a natural choice as a framework for algorithm configuration procedures. Indeed, Graham and Leyton-Brown (2024) argue that UCB is the best way to apply utilitarian algorithm configuration to infinite parameter spaces, and show that using UCB is a significant improvement over configuration procedures that implement other bandit procedures. However, UCB is essentially designed for regret minimization, while the algorithm configuration problem is a case of best arm identification. Thankfully, the LUCB (Lower Upper Confidence Bound) algorithm (Gabillon, Ghavamzadeh, and Lazaric 2012; Kalyanakrishnan et al. 2012) is a variant of UCB that was designed to be applied to the best arm identification problem. Actually, the LUCB algorithm was designed to solve the $m$-best arm identification problem, where a learner is tasked with identifying the top $m$ arms from a set of $n$ arms. Although we are only concerned with the special case of $m=1$, it is easier to understand what LUCB is doing by considering the more general case. In each iteration, LUCB selects and samples two arms (i.e., configurations). First it samples the arm with the smallest lower confidence bound from the $m$ empirically best arms. Next it samples the arm with the largest upper confidence bound from among the nmn-m remaining arms. This way, the LUCB algorithm attempts to tease apart the $m$ best arms from the nmn-m worst arms. Figure 2 (Left) in Section 3 shows an example of the improvement achieved by using the LUCB algorithm. This is drastic early in a search.
@@ 2.3 Adding New Configurations
The second condition does have this property. At any point during its execution, COUP will recommend the configuration $i^{*}=\arg\max LCB_{i}$, the configuration with largest lower confidence bound. This is the configuration about which it can prove the best guarantee. The quantity $LCB_{i^{*}}$ is the highest expected utility that COUP can guarantee at any time. We would like to push $LCB_{i^{*}}$ as close to 1 as possible because then we will have (i) a very good configuration, and (ii) a very strong guarantee about this configuration. So we can think of the quantity $1-LCB_{i^{*}}$ as the utility “still on the table”; the largest amount of utility by which we might hope to improve our guarantee. We can write $1-LCB_{i^{*}}=1-UCB_{i^{\prime}}+UCB_{i^{\prime}}-LCB_{i^{*}}$. The term $UCB_{i^{\prime}}-LCB_{i^{*}}$ is the best $\epsilon$ we can guarantee with respect to our existing set of configurations, the maximum amount by which we can shrink $1-LCB_{i^{*}}$ by running existing configurations. On the other hand, since $UCB_{i^{\prime}}$ is the largest upper bound from our current sample of configurations, we can be confident that the configurations we have seen so far have utility at most $UCB_{i^{\prime}}$. So the term $1-UCB_{i^{\prime}}$ is the utility potentially to be gained from other configurations that we have not yet seen. This is illustrated in Figure 1.
@@ 2.4 Model-guided Configuration Search
COUP can be improved by incorporating a learning model that predicts the performance of candidate configurations before spending significant amounts of time investigating them. If good configurations are rare enough, then we will be all but guaranteed to never encounter them by random sampling alone. A model will at least have the possibility of finding such configurations, and there is evidence for this: Hutter, Hoos, and Leyton-Brown (2011) report that SMAC outperforms ROAR, its model-free variant, in all scenarios and substantially outperforms ROAR on some large configuration spaces, indicating that at least in some cases the prediction model is able to find significantly better configurations than random sampling alone.
As COUP runs, it generates the performance data needed to train a performance model and we can use the trained model to help identify promising areas of the search space to investigate. Areas that are predicted to yield configurations with high performance will be investigated more, as will areas that have high uncertainty. We use a gradient-boosted set of regression trees as our base learning model to predict configuration performance. We use the XGBoost (Chen and Guestrin 2016) implementation for gradient boosting. In order to produce uncertainty estimates we perform a bootstrap aggregation procedure, sampling subsets of configurations and training an XGBoost model on each one. For each candidate configuration we make a prediction with each of the models, giving us a distribution of values. We then consider 95% confidence intervals, giving us an uncertainty band around each prediction. For more details please see Appendix A.
Given a trained model, we employ the same procedure used by SMAC (Hutter, Hoos, and Leyton-Brown 2011) to identify promising configurations to investigate. Briefly, we start with the ten best configurations seen so far and perform a local search starting with each. For discrete parameters, we consider the set of all configurations that differ in exactly one parameter value. For continuous parameters, we normalize values to the range $[0,1]$ and sample four neighbors according to a truncated Gaussian distribution with mean equal to the parameter value and standard deviation equal to 0.2. We continue this search as long as the neighbors have a larger upper confidence bound. Once we have sampled 10 neighbors, we use the model to predict their performance, as well as the performance of 10,000 randomly selected configurations, choosing the one that has the highest predicted 95% upper confidence bound as the next candidate.

### v2

@@ Abstract
Utilitarian algorithm configuration identifies a parameter setting for a given algorithm that maximizes a user’s utility. Utility functions offer a theoretically well-grounded approach to optimizing decision-making under uncertainty and are flexible enough to capture a user’s preferences over algorithm runtimes (e.g., they can describe a sharp cutoff after which a solution is no longer required, a per-hour cost for compute, or diminishing returns from algorithms that take longer to run). COUP is a recently-introduced utilitarian algorithm configuration procedure which was designed mainly to offer strong theoretical guarantees about the quality of the configuration it returns, with less attention paid to its practical performance. This paper closes that gap, bringing theoretically-grounded, utilitarian algorithm configuration to the point where it is competitive with widely used, heuristic configuration procedures that offer no performance guarantees. We present a series of improvements to COUP that improve its empirical performance without degrading its theoretical guarantees and demonstrate their benefit experimentally. Using a case study, we also illustrate ways of exploring the robustness of a given solution to the algorithm selection problem to variations in the utility function.
Codehttps://github.com/drgrhm/practical-coup
@@ 1 Introduction
Algorithm configuration is the process of finding a complete setting of an algorithm’s parameters that lead it to perform well on a given set of input instances. Often, the algorithm of interest is a solver for a hard computational problem such as Boolean Satisfiability (SAT). Since SAT is NP-complete, we are unlikely to be able to find a single solver or configuration that performs well on all possible inputs. In practice, however, we are never interested in solving all possible inputs. Instead, we tend to be interested in solving input instances drawn from a given distribution.
Algorithm configuration is thus a bandit problem (lattimore2020bandit) where each configuration is an arm, and running a configuration on an input instance means drawing a sample from that arm’s underlying distribution. Unlike the standard bandit setting, in which there is a unit cost associated with pulling an arm, in algorithm configuration the cost is the time it takes for the algorithm to solve the input. We can also choose whether each such evaluation should be censored (“capped”) at some maximum runtime. The objective in algorithm configuration is not regret minimization, but rather (approximate) best arm identification.
@@
Algorithm configuration procedures have traditionally aimed to minimize the expected runtime of the returned configuration. The most practical such methods make these decisions according to carefully chosen heuristics (birattari2002racing; hutter2009paramils; ansotegui2009gender; hutter2011sequential; lopez2016irace; SMAC3). A second line of work has identified (considerably less practical) algorithms that offer worst-case guarantees about the runtime required to identify approximately optimal solutions (kleinberg2017efficiency; weisz2018leapsandbounds; kleinberg2019procrastinating; weisz2019capsandruns; weisz2020impatientcapsandruns). For the purposes of this paper, we define an optimality guarantee as depending on three parameters: $\epsilon$, $\gamma$, and $\delta$. With failure probability $\delta$, we want to prove both that the returned configuration is within an additive factor of $\epsilon$ from the best configuration sampled, and that the probability of sampling a configuration that is better than the best one sampled so far is at most $\gamma$.
@@
It has recently been argued that algorithm configuration should rest on the firmer decision-theoretic foundation of maximizing a utility function over algorithm runtime (graham2023formalizing). That is, let $u$ be a function that maps runtimes to utility values in $[0,1]$, encoding the user’s preferences about runtimes of different lengths; we should then maximize $u$’s expected value. A recent line of work has proposed procedures that offer worst-case guarantees about the runtime required to approximately optimize a utilitarian objective (graham2023utilitarian; brandt2023ac). The most recent such procedure is COUP (Continuous, Optimistic Utilitarian Procrastination; graham2024utilitarian), which was shown empirically to perform favorably with respect to other configuration procedures that offer (in some cases, somewhat incomparable) optimality guarantees: Successive Halving (jamieson2016non), Hyperband (li2018hyperband), AC-Band (brandt2023ac), ImpatientCapsAndRuns (weisz2020impatientcapsandruns) and Structured Procrastination with Confidence (kleinberg2019procrastinating). Pseudocode for COUP is included in Appendix A, but in essence it is an instantiation of the UCB algorithm (lai1985asymptotically; lai1987adaptive), coupled with a captime exploration technique and a condition for expanding the set of configurations under consideration. For each configuration $i$, COUP maintains upper and lower confidence bounds $UCB_{i}$ and $LCB_{i}$. In each iteration it chooses the configuration $i^{\prime}=\arg\max_{i}UCB_{i}$ and performs a run of it. In this way it works to improve $\epsilon^{*}=\max_{i}UCB_{i}-\max_{i}LCB_{i}$, which is the best optimality guarantee it can make. If COUP detects that the captime it is using is too small, it will double it. If COUP detects that its set of configurations is too small, it will add more. COUP is anytime in its guarantees with respect to both $\epsilon$ and $\gamma$: as it runs it makes tighter and tighter optimality guarantees with respect to larger and larger sets of configurations. COUP is also adaptive to the inputs it is given: it provably makes better guarantees in less time when given easier inputs.
@@
This paper begins in Section 2 by describing a series of improvements to COUP that boost its empirical performance without compromising its theoretical guarantees. First, we refine the confidence bounds used by COUP, making the optimization process more efficient. Second, we incorporate a different version of the UCB algorithm that specifically targets the best arm identification problem. Third, we suggest a new condition for when COUP should add new configurations, which frees the user from having to provide a set of parameters ahead of time, instead adapting to the inputs it sees. And fourth, we make COUP model-based, using a boosted forest of regression trees to help COUP find new configurations. In Section 3 we show for the first time that an algorithm configuration method offering theoretical guarantees (our modified version of COUP) robustly achieves performance roughly equivalent to that of SMAC (hutter2011sequential; SMAC3), a heuristic procedure widely seen as the empirical state of the art. Finally, in Section 4, we address a shortcoming of the utilitarian framework: that it is not always easy for the end user to choose the exact utility function to optimize. Via a case study on data from one of the most prominent international evaluations of heuristic algorithms, the SAT Competition, we illustrate how runtime data can be used to reason about the utility functions for which a given algorithm would be considered optimal.
@@ 2.1 Improved Confidence Bounds
A key technical tool used in COUP is Hoeffding’s Inequality (hoeffding1963pobability). This concentration inequality asserts that for bounded random variables, the probability that the observed sample mean is far from the true population mean is exponentially small in the number of samples taken. However, Hoeffding’s Inequality involves a key approximation, which is tight when the true mean is close to $1/2$ (in the case of random variables bounded on [0,1]), but grows worse when the true mean is close to either 0 or 1—a situation encountered often in utilitarian algorithm configuration. We can do away with this approximation by appealing directly to the “Chernoff–Hoeffding Lemma”, and solving numerically for the confidence bounds, which are then guaranteed to be tighter than a naive application of Hoeffding’s Inequality; for a detailed exposition see lattimore2020bandit, Chapter 10. The confidence bounds come from solving an optimization problem that bounds the KL-Divergence between the observed sample mean and the bounding quantity. We thus refer to them as “KL bounds”.
COUP maintains a captime $\kappa$, which is the maximum time it will spend on each run of a configuration. To derive upper and lower confidence bounds for a configuration $i$’s mean utility, COUP needs to estimate its capped mean runtime $U_{i}=\mathbb{E}_{t_{i}}[u(\min(t_{i},\kappa))]$ and its probability of not capping $F_{i}=\Pr_{t_{i}}(t_{i}\leq\kappa)$. Both of these estimates can be improved by switching from Hoeffding bounds to KL bounds. Define $d(p,q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$. Suppose $n$ is the current number of configurations, $m$ is the number of samples we have taken for configuration $i$, and $\kappa$ is the captime we have taken them at. If $\widehat{U}_{i}$ is the empirical capped average runtime, $\widehat{F}_{i}$ is the empirical fraction of completed runs, and $a=\frac{1}{m}\ln\frac{27n^{2}m^{2}(\log(\kappa+1)^{2}}{\delta}$ then
@@
are confidence bounds that hold for all configurations at all times during COUP’s execution with probability at least $1-\delta$ (detailed derivations can be found in Appendix B). This alteration essentially makes COUP an implementation of the KL-UCB algorithm (cappe2013kullback) for bandit problems. In Figure 2 (Left) in Section 3 we show that COUP performs better when using these refined bounds.
@@ 2.2 Lower Upper Confidence Bound
The UCB algorithm is a natural choice as a framework for algorithm configuration procedures. Indeed, graham2024utilitarian argue that UCB is the best way to apply utilitarian algorithm configuration to infinite parameter spaces, and show that using UCB is a significant improvement over configuration procedures that implement other bandit procedures. However, UCB is essentially designed for regret minimization, while the algorithm configuration problem is a case of best arm identification. Thankfully, the LUCB (Lower Upper Confidence Bound) algorithm (gabillon2012best; kalyanakrishnan2012pac) is a variant of UCB that was designed to be applied to the best arm identification problem. Actually, the LUCB algorithm was designed to solve the $m$-best arm identification problem, where a learner is tasked with identifying the top $m$ arms from a set of $n$ arms. Although we are only concerned with the special case of $m=1$, it is easier to understand what LUCB is doing by considering the more general case. In each iteration, LUCB selects and samples two arms (i.e., configurations). First it samples the arm with the smallest lower confidence bound from the $m$ empirically best arms. Next it samples the arm with the largest upper confidence bound from among the nmn-m remaining arms. This way, the LUCB algorithm attempts to tease apart the $m$ best arms from the nmn-m worst arms. Figure 2 (Left) in Section 3 shows an example of the improvement achieved by using the LUCB algorithm. This is drastic early in a search.
@@ 2.3 Adding New Configurations
The second condition does have this property. At any point during its execution, COUP will recommend the configuration $i^{*}=\arg\max LCB_{i}$, the configuration with largest lower confidence bound. This is the configuration about which it can prove the best guarantee. The quantity $LCB_{i^{*}}$ is the highest expected utility that COUP can guarantee at any time. We would like to push $LCB_{i^{*}}$ as close to 1 as possible because then we will have (i) a very good configuration, and (ii) a very strong guarantee about this configuration. So we can think of the quantity $1-LCB_{i^{*}}$ as the utility “still on the table”; the largest amount of utility by which we might hope to improve our guarantee. We can write $1-LCB_{i^{*}}=1-UCB_{i^{\prime}}+UCB_{i^{\prime}}-LCB_{i^{*}}$. The term $UCB_{i^{\prime}}-LCB_{i^{*}}$ is the best $\epsilon$ we can guarantee with respect to our existing set of configurations; the maximum amount by which we can shrink $1-LCB_{i^{*}}$ by running existing configurations. On the other hand, since $UCB_{i^{\prime}}$ is the largest upper bound from our current sample of configurations, we can be confident that the configurations we have seen so far have utility at most $UCB_{i^{\prime}}$. So the term $1-UCB_{i^{\prime}}$ is the utility potentially to be gained from other configurations that we have not yet seen. This is illustrated in Figure 1.
@@ 2.4 Model-guided Configuration Search
COUP can be improved by incorporating a learning model that predicts the performance of candidate configurations before spending significant amounts of time investigating them. If good configurations are rare enough, then we will be all but guaranteed to never encounter them by random sampling alone. A model will at least have the possibility of finding such configurations, and there is evidence for this: hutter2011sequential report that SMAC outperforms ROAR, its model-free variant, in all scenarios and substantially outperforms ROAR on some large configuration spaces, indicating that at least in some cases the prediction model is able to find significantly better configurations than random sampling alone.
As COUP runs, it generates the performance data needed to train a performance model and we can use the trained model to help identify promising areas of the search space to investigate. Areas that are predicted to yield configurations with high performance will be investigated more, as will areas that have high uncertainty. We use a gradient-boosted set of regression trees as our base learning model to predict configuration performance. We use the XGBoost (chen2016xgboost) implementation for gradient boosting. In order to produce uncertainty estimates we perform a bootstrap aggregation procedure, sampling subsets of configurations and training an XGBoost model on each one. For each candidate configuration we make a prediction with each of the models, giving us a distribution of values. We then consider 95% confidence intervals, giving us an uncertainty band around each prediction. See Appendix E for more details.
Given a trained model, we employ the same procedure used by SMAC (hutter2011sequential) to identify promising configurations to investigate. Briefly, we start with the ten best configurations seen so far and perform a local search starting with each. For discrete parameters, we consider the set of all configurations that differ in exactly one parameter value. For continuous parameters, we normalize values to the range $[0,1]$ and sample four neighbors according to a truncated Gaussian distribution with mean equal to the parameter value and standard deviation equal to 0.2. We continue this search as long as the neighbors have a larger upper confidence bound. Once we have sampled 10 neighbors, we use the model to predict their performance, as well as the performance of 10,000 randomly selected configurations, choosing the one that has the highest predicted 95% upper confidence bound as the next candidate.